# AI-Driven Diabetic Retinopathy Screening: Multicentric Validation of AIDRSS in India

### v1

@@ 2.1 Study Design
This study employed a month-long observational design to document demographic information and evaluate the efficacy of an Artificial Intelligence-based Diabetic Retinopathy Screening System (AIDRSS) in a low socio-economic status (SES) community. The study included capturing retinal fundus photographs using a portable non-mydriatic automatic fundus camera supported by the AIDRSS for image analysis.
@@ 2.5 Proposed Algorithm
Deep learning is currently state-of-the-art for computer vision/image processing, speech, text problems, and automotive. The AIDRSS utilized a deep learning algorithm with a modified neural network architecture comprising approximately 50 million trainable parameters and 250 layers deep. The proposed network is a modified architecture version presented in [6]. The values of the network parameters are learned using the Adam optimizer algorithm. Neural network hyper-parameter learning is an iterative process based on loss value computed between ground truths given by ophthalmologists and network predictions. This loss value acts as a feedback to the optimizer to learn the most representative network parameters to understand retinal pathological lesions.
@@ 5 Conclusion
Informed consent was obtained from all participants for capturing their fundus image using the CrystalVue NFC600 device and for the prospective study of their data for synthetic image generation.

### v2

@@ 2.1 Study Design
This was a retrospective study using data collected from routine screenings at multiple centers in Kolkata, India. The data comprised retinal fundus images collected between January and March 2023. A total of 5,029 participants were screened, and their de-identified retinal images were analyzed retrospectively. The validation was performed on a withheld dataset from a separate site, ensuring an independent evaluation of AIDRSS.
2.2 Validation Methodology
We validated the AIDRSS system’s performance on a completely withheld dataset from an independent screening site, comprising 1,200 images not included in the training or development datasets. Additionally, five-fold cross-validation was employed to assess the model’s generalizability across diverse patient subsets. Each fold reserved data from a unique center, ensuring no overlap between training and testing images.
The withheld site validation tested the system’s robustness when applied to a new population, while cross-validation provided insight into its consistency across different subsets of the study population. Performance metrics, including sensitivity, specificity, positive predictive value (PPV), and negative predictive value (NPV), were calculated for both the cross-validation and withheld datasets.
To ensure statistical robustness, confidence intervals (95% CI) for all performance metrics were computed using bootstrap resampling with 1,000 iterations. This method provides a non-parametric estimate of the variability in the calculated metrics, offering a reliable measure of the system’s performance. Additionally, McNemar’s test was performed to compare the performance of the AIDRSS against baseline or other models on paired data. These statistical approaches ensure a thorough and accurate assessment of the model’s diagnostic capabilities.
@@ 2.6 Proposed Algorithm
Deep learning is currently state-of-the-art for computer vision/image processing, speech, text problems, and automotive. The AIDRSS utilized a deep learning algorithm with a modified neural network architecture comprising approximately 50 million trainable parameters and 250 layers deep. The proposed network is a modified architecture version presented in [6]. The values of the network parameters are learned using the Adam optimizer algorithm. Neural network hyper-parameter learning is an iterative process based on the loss value computed between ground truths given by ophthalmologists and network predictions. This loss value acts as a feedback to the optimizer to learn the most representative network parameters to understand retinal pathological lesions.
@@ 3.3 Analysis of Referable DR Detection
3.4 Comparison with Established AI Models
To contextualize the performance of AIDRSS, we compared it with established AI models for diabetic retinopathy screening, specifically IDx-DR [9], EyeArt [10], and AEYE-DS [11]. These models have been previously validated and are widely recognized in the field. Table 3 summarizes the sensitivity and specificity metrics for AIDRSS and the benchmark systems.
Table 3: Performance Metrics Comparison with Established AI Models
Metric AIDRSS IDx-DR EyeArt AEYE-DS
Sensitivity (%) 92.0 87.4 96.0 93.0
Specificity (%) 88.0 89.5 88.0 91.4
These results demonstrate that AIDRSS performs competitively with other state-of-the-art systems. Notably, AIDRSS achieves a balance between sensitivity and specificity, which is particularly advantageous in resource-limited settings. High sensitivity ensures that most cases of diabetic retinopathy (DR) are detected, reducing the risk of missed diagnoses and enabling early intervention. At the same time, high specificity minimizes false positives, thereby preventing unnecessary referrals and reducing the burden on limited healthcare resources. This balance enhances the efficiency and effectiveness of screening programs, making AIDRSS a valuable tool in areas with constrained healthcare infrastructure.
However, it is important to consider the limitations of comparing AIDRSS with other AI models. Differences in datasets, imaging equipment, and population characteristics can significantly influence performance metrics. For instance, variations in image quality, camera types, and demographic factors such as ethnicity and age may affect sensitivity and specificity outcomes. AI models trained on datasets from one region may underperform when validated on external datasets due to differences in disease prevalence and imaging conditions. These factors highlight the importance of accounting for dataset composition and generalizability when interpreting comparative performance metrics.
Future work should focus on validating AIDRSS in diverse geographic and clinical settings to assess its robustness across different populations. Such efforts will further establish its utility as a scalable, reliable solution for diabetic retinopathy screening.
@@ 5 Conclusion
Informed consent was obtained from all participants for capturing their fundus image using the CrystalVue NFC600 device. Participants were informed that their de-identified data might be used for further research and analysis.